OpenStudy (thecatman):

Which literal equation is equivalent to b equals five times a minus four? a equals four times b plus five a equals negative five plus four times b a equals b plus four all over five a equals b minus four all over five

OpenStudy (thecatman):

@Xmoses1

OpenStudy (thecatman):

please help

OpenStudy (***[isuru]***):

first convert it in to an equation with symbols.. b= 5a - 4 now solve it for "a" b + 4 = 5a -4 + 4 5a = b + 4 a = (b + 4) / 5 which is a equals b plus four all over five